A GRAPHICAL-USER-INTERFACE FOR THE ASSIGNMENT OF ROTATIONALLY RESOLVED SPEC-TRA

\underline{JOCHEN KÜPPER}, (current address: University of North Carolina, Department of Chemistry, Chapel Hill, NC 27599) Heinrich-Heine-Universität, Institut für Physikalische Chemie, 40225 Düsseldorf, Germany.

The program krot for the simulation, assignment and fitting of rotationally resolved spectra is presented. A graphical user interface has been developed based on the KDE^{a} libraries. It displays several experimental and simulated spectra simultaneously and allows one to make assignments using a mouse-driven interface. Currently the program uses a single engine for the calculation and least-squares-fitting of spectra, but provisions are made to call any program or set of routines to do the actual calculations. The current calculation program arnirot^{b} allows the calculation of microwave and vibrational/vibronic spectra of asymmetric rotors, including centrifugal distortion terms in Watson’s A- or S-reduction. Moreover linear terms for the description of axis reorientation or coupling to internal motions are included. A link to the download page is given at http://www-public.rz.uni-duesseldorf.de/˜ mschmitt/HRLIF.htm.

aThese libraries are OpenSource and available at http://www.kde.org.

bThis program is based on the program blzyrot, Blaise B. Champaigne and David W. Pratt, private communication, 1994.